Determine the general solution to the equation ∂2u/∂t2=c2(∂2u/∂x2) under the boundary conditions u(0,t)=u(1,t)=0 and initial conditions u(x,0)=Φ(x), ut(x,o)=Ψ(x)


1
Expert's answer
2021-11-18T12:57:40-0500

solution of the one-dimensional wave equation:

"u(x,t)=\sum A_nsin(\pi nx/l)cos(\pi nct/l)"

where

"A_n=\frac{2}{l}\int^l_0 u(x,0)sin(\pi nx/l)dx=\frac{2}{l}\int^l_0 \Phi (x)sin(\pi nx/l)dx"


then:

"u_t(x,t)=-\frac{A_n\pi c}{l}sin(\pi x/l)sin(\pi ct/l)"

"u_t(x,0)=\u03a8(x)=0"


l is length of object where wave occurs (for example, string)